OpenStudy (tylermckinney16):

Help please. Write the equation of a parabola with vertex at the origin and directrix y=-6. Show your steps.

OpenStudy (tylermckinney16):

@Aveline @IconForHire56 @knesha

OpenStudy (aveline):

The standard equation of a parabola is: \[(x-h)^{2}=4p \times (y-k)^{2}\] The directrix is: \[y=k-p\] And the focus is: \[(h, k + p)\] (h,k) is the vertex

OpenStudy (aveline):

\[(x – h)^{2}= 4p \times (y – k) \] Let's plug in (0,0) for the vertex since the vertex is at the origin: \[x ^{2}=4p \times y\] The directrix is -6 units away from the vertex, so p=6 \[x ^{2}=24y\]

OpenStudy (aveline):

I noticed I made an error on my first post: The standard equation is \[(x-h)^{2}=4p \times (y-k)\]

OpenStudy (tylermckinney16):

Ok thanks .

OpenStudy (welshfella):

sorry i got called away in any case you had great help

OpenStudy (welshfella):

the parabola will open upwards with the focus at y=6

OpenStudy (tylermckinney16):

ok.

OpenStudy (welshfella):

https://www.desmos.com/calculator/6brcymnbly there it is

OpenStudy (tylermckinney16):

Is X^2=24Y the answer?